### v1

# Large earthquakes follow highly unequal ones

Abstract
It was conjectured for a long time that the tectonic plates are in a self-organized state of criticality and that the Gutenberg-Richter (power) law is a manifestation of that. It was recently shown that for a system near criticality, the inequality of their responses toward external driving could indicate proximity to the critical point. In this work, we show with numerical simulations and seismic data analysis that large earthquake events have a tendency to follow events that are highly unequal. We have applied this framework to various tectonically active regions, such as North America, Southern Japan, parts of South-East Asia and Indonesia.
I Introduction
When tectonic plates move (at a very slow rate), they undergo a stick-slip motion that causes earthquakes when the stored potential energy due to the motion is released due to a sudden slip event. While much is known through more than a century of studies related to the motions and the subsequent energy released, its understanding still faces severe difficulties, particularly due to the inaccessibility of the regions of origin (often tens of kilometers underground) [1]. However, the phenomenon of earthquake has fascinated researchers from various disciplines alike not only due to their devastating consequences, but also due to their remarkable regularities in the statistical sense [2].
Particularly, it is known that the earthquake magnitudes follow the Gutenberg-RIchter law [3], $\log_{10}N=a-bM$, stating that the number of earthquakes ($N$) of magnitude $M$ are related exponentially ($a$ and $b$ are constants). It is also observed that the rate of aftershocks decays with time in a power-law $n(t)=K/(t+c)^{p}$, known as the Omori-Utsu law [4, 5]. There are other empirical laws about earthquake dynamics as well, which are verified using data analysis across various regions in the world. While a complete understanding of the mechanism behind the earthquake phenomenon is still lacking, the empirical observations have been useful in gaining insights into the earthquake dynamics, improving the prospective to model the phenomenon and to also aim for a forecasting possibility, which has so far largely evaded understanding.
There have been some efforts in searching for regularities in earthquake statistics that could potentially lead to hazard analysis. One such area is the change in the exponent value prior to a large event [6, 7, 8]. This methods follows from the stress dependence of avalanche size exponent in sheared rocks, and also in fracture. In such cases, the avalanche size exponent value decreases prior to a large event. In earthquake also, when expressed in terms of energy, the size distributions follow a power law. The values of the power law exponent decreases, if the events are sampled prior to a large event.
In this work, we analyze a regularity, which depends on the inequality of the earthquake energy releases. Particularly, we note that prior to a large earthquake, the energy released are more unequal than when such large events are not happening. Inequality in responses of a driven disordered SOC-like systems is known to be useful in terms of its universal features [9, 10]. It has also shown promises, both in models and (compression) fracture experiments, to have forecasting capabilities of imminent large events [11, 12]. Here we have measured inequality of energy released within a time-window for earthquake time series in various models and also in the data in several regions. It has then been shown to have a positive correlation with ”large” events that immediately follow such a time-window. We have quantified the earthquake inequalities through Gini and Kolkata indices, and then shown its positive correlation with the events that follow.
For the remaining part of the paper, we first describe the models used and the method of measuring the inequalities in their responses. We then correlate the inequalities of the responses with subsequent events. Finally, we compare with data from different regions on earth and show that qualitatively such correlations hold, indicating an underlying principle of relating response inequalities with proximity to a critical point.
II Models and methods
Here we first describe the models used for this study and then move on to define the inequality indices that are used.
II.1 Train model of earthquake
The train model of earthquake is a simplification on the Burridge-Knopoff model [13], where the stick-slip motion of the overlapping tectonic plates are modeled through massive blocks, connected through linear elastic springs, and driven over a rough surface. The frictional force creates the stick-slip motion and the (elastic) energy released can be measured through the motion of the blocks.
We consider a further simplification on the model [14], where we take the frictional surface to be discrete set of blocks, but fixed in space. When the “train” moves over it, slowly pulled by the rightmost block (“engine”), say, some blocks of the train overlaps with the fixed blocks below. At that point, they have a random pinning force that oppose the motion, and is drawn from a uniform distribution in (0,1) randomly. The forces on a block in the train at any given instance are, therefore, the elastic forces from its nearest neighbors due to potential stretching, and the random pinning force. If the pull due to the elastic force exceeds the random pinning, then the block is moved towards the direction of the pulling, hopping over to the next block below, only if that space is not occupied already. The train only moves along the direction of pulling and never opposite.
The dynamics then proceeds as follows, the right most block is moved further to the right by one step and the forces are checked on all the other blocks. Some blocks might move and that can in-turn result in movements of some other blocks and so on. The rightmost block is held fixed during this re-adjustments. When all blocks stop moving, the rightmost block is moved again one step and so on. The number of blocks moving between two successive movements of the “engine” is called an avalanche. The self-organized critical dynamics of this model is well studied and can be shown to be equivalent to the quenched Edwards-Wilkinson model [15]. The avalanche size distribution follow a power-law in one and two dimensions. Here we only study the one dimensional version of the model.
II.2 Sandpile-like model of earthquake
We adopt the model introduced in ref. [16] for this. Sandpile models typically measure event size by the number of topplings, the present model defines physically meaningful earthquake observables. As a result, a Gutenberg–Richter (Power) law emerges only in a specific parameter regime. We consider a two-dimensional square lattice consisting of $N$ sites. The stress at each site increases uniformly and linearly in time at a constant rate ($v_{0}$). When the stress at a particular site reaches a given threshold value ($S_{c}$), an earthquake event is triggered. We choose $D_{1}$ as a constant stress threshold. Sites with stress greater than $S_{c}-D_{1}$ are identified and grouped into clusters of nearest neighbors. Only the cluster containing the triggering site participates in the earthquake. For every site in this cluster, reduce its stress to a new stress value $S_{c}-D_{s}$, where $D_{s}=D_{2}+D_{3}$. Here, $D_{2}$ is a fixed paramater and $D_{3}$ is a random parameter. In this model ”Cluster” defines a set of connected sites (grids) where stress level exceed the choosen threshold and ”Avalanche” defines as a specific cluster which contains the site with maximum stress (epicenter) at that time. If $n$ is the number of sites in this cluster then, the resulting seismic moment ($M$) and earthquake magnitude ($m$) are related as,
$m=\log M=\frac{3}{2}\log n$
[…]
$D_{3}\ll D_{1}\ll D_{2}$
which means that, earthquakes involve a stress release, but the variations in stress drop are small relative to the stress level at which events are initiated. We introduce another spatial correlation parameter ($D_{d}$), where the tresholds are correlated within a typical size of $D_{d}^{2}$ (see e.g., [17]). Large $D_{d}$ value will give large avalanche size.
II.3 Inequality indices used
In quantifying the inequality of the earthquake events, in simulations and in the data, we use inequality indices that are traditionally used to quantify inequality of wealth distribution in a population, viz. the Gini [18] and Kolkata [19] indices. In measuring that we need to first construct the Lorenz curve. It is done as follows: For a set of numbers, here representing earthquake energy released, are first sorted in ascending order. Then $L(p)$ is measured, which is the cumulative fraction of the smallest $p$ fraction of the events. Clearly, $L(0)=0$ and $L(1)=1$. This is called the Lorenz curve. If all events were of exactly equal size, then this would be a diagonal line from (0,0) to (1,1). Any departure from the diagonal line is, therefore, a measure of inequality of the events. The Gini index is defined as the area between the diagonal line and the Lorenz curve, normalized by the area under the diagonal line (1/2). Formally, $g=1-2\int\limits_{0}^{1}L(p)dp$. The value of $g$ is always between 0 and 1, representing complete equality and complete inequality, respectively.
The Kolkata index is defined as the largest 1k1-k fraction of the events that accounts for the $k$ fraction of the cumulative energy released in this case. Formally, it can be obtained by solving 1k=L(k)1-k=L(k). It is a generalization of the Pareto’s 80-20 law in wealth (and many other) distributions, that states that 80% of wealth is typically possessed by the 20% of the richest individuals. Here it gets modified as 1k1-k fraction of the largest events accounts for the $k$ fraction of the total energy released.
These measured have been used recently in responses in systems approaching criticality and have been found to have universal features (see e.g., [20]). Here we use these for earthquake statistics (see also [21]).
III Results
It is known that in systems approaching a critical point, the responses tend to approach a diverging point and the corresponding inequality measures $g$ and $k$ becomes nearly equal before the system reaches the critical point [20, 10]. For Self Organized Critical (SOC) systems also, this near universal behavior is seen [9].
Figure 1: The top figure shows the time series plot of the avalanches for the train model. The bottom figure is the same series, but now sorted according to the corresponding Gini index associated with each event (calculated from the immediately preceeding 100 events. The larger avalanches tend to accumate near larger $g$ values.).
Figure 2: The binned avalanches are shown as a function of $g$ values, with a binning range of 0.01 in the $g$ values for the different system sizes in the 1-d train model. A clear postive correlation is observed.
First we apply this to the response statistics of the train model. We choose a window of 100 consecutive events from the time series of the avalanche sizes (mimicking energy release here) and calculate $g$ and $k$. We then assign that $g$ and $k$ values to the event immediately following the 100 event window. We go on to do this assignments as we slide the window in time by one event each time. This means that all events, except the first 100, have an inequality index value assigned to it. Then we take all those labelled events and sort them in ascending order of the $g$ (and $k$) values. Fig. 1 shows the unsorted and the sorted time series. There is a tendency of accumulation of the larger events towards larger $g$ value. A similar trend (not shown) is observed when the sorting is done with respect to the $k$ values.
Figure 3: The top figure shows the time series of the avalanches seen in the 2d SOC model. The bottom figure is the same time series, but now sorted according to the $g$ value associated with each event (calculated from the immediately preceeding 100 events). Similar to the train model, a clear correspondence is noted with high $g$ and large avalanches, except for the vary high $g$ values, where events are less.
Figure 4: The binned avalanches with $g$ values for the 2d SOC model. The bin size is 0.01. A positive correlation with increasing $g$ is observed, as in the case of the train model.
We then also plot, in Fig. 2, the total energy released for a particular $g$ values (binned in an interval of 0.01). It shows a clear increase with $g$, before it suddenly drops for very high $g$ values, where events are very small in number.
Figure 5: Earthquake energy distribution plots of real earthquake data. Earthquake magnitude was converted into energy using the relation $E\propto 10^{1.5M}$, where $M$ denotes the earthquake magnitude. The resulting energy is expressed in joules. This formula is taken from Peter Bormann et al. [23], in this paper they proposed $E=10^{1.5M}+4.8$, but in our method we ignore the constant term as it do not effect the $g$ value, only the magnitude of earthquake energy increases. Exponent value for each region is mention on the right-most figure. The common fitted Power law exponent is $b_{c}=-1.733\pm 0.002$ with KS D (Kolmogorov-Smirnov Distance)=0.166=0.166.
Figure 6: Top row represents earthquake energy in Joule vs Year for different regions, to make these plots we take every 100100 events. Bottom row represents the same events but now sorted according to their $g$ values. The earthquake data are taken from USGS earthquake catalog from year 197520251975-2025. We calculate gkg-k index by using sliding window approach. We convert the ISO time scale into UTC time scale and then apply different window sizes $W=25,50,100,200,400$ and $800$ to calculate gini and Kolkata inequality indices. The calculated gkg-k is assigned to the very next event which is outside the window. These plots show similar trend as those in the model.
Next we do the same exercise for the 2d sandpile like model. In Fig. 3, the unsorted (original time ordered) and the sorted (according to $g$ values) event series are shown. A clear trend of accumulation of larger events towards higher $g$ value is again observed. The plot (Fig. 4) of the energy released within a fixed interval of $g$ values show a peak, typically around 0.87, which is also known to be the near universal trend in SOC models. The curve then drops due to lack of events with very high $g$ values.
Finally, we apply this to the real data of earthquake catalogs in different parts of the world (Southern Japan, South-East Asia, North America and Indonesia). The coordinates used are as follows: Southern part of Japan: Minimum latitude 30.145, Maximum latitude 41.278 Min longitude 129.792, Max longitude 142.119. South East Asia: Min latitude 20.633, Max latitude 35.747 Min longitude 67.588, Max longitude 99 Western part of North America: Min latitude 12.383, Max latitude 61.143 Min longitude (-145.547), Max longitude (-46.318) Indonesia: Min latitude (-7.537), Max latitude 5.703 Min longitude 95.142, Max longitude 107.051. The catalog used (USGS) can be found in [22].
First we check that the data set show Gutenberg-Richter scaling (see Fig. 5). The lower cut-off used in magnitude is 4.5. Then the magnitudes are conveted to energy [23] (inequalities in magnitudes were reported very recently in [21]). In Fig. 6, the unsorted (actual order) and sorted (according to $g$ values) can be seen. Once again, a clear trend of accumulations of larger events towards higher $g$ values are observed. In Fig. 7, the energy released within a fixed range (of interval 0.01) $g$ values are shown. It also shows a clearly increasing trend, as in the case of the models mentioned above.
IV Discussions and conclusion
The inequalities of the earthquake energy released, in models and data, show a consistent trend of the larger events following a set of highly unequal ones. This is seen in the two models studied here (the train model of earthquake [14] and the 2d SOC like model [16]) and also in data taken from various earthquake prone zones around the world. While earthquake have been cast in the form of SOC dynamics over many years, there are some persistent doubts as well. However, at least for the behavior of inequality indices this seems to follow the near universal trend observed in the SOC models [9].
Figure 7: These plots represent the total earthquake energy in every $0.01$ step of Gini value. We also indicate the $g_{max}$ which represents the $g$ values of the maximum total earthquake energy.
In the context of earthqaukes, this also aligns with the earlier observation of lowering of the size distribution exponent prior to large events [7]. A lowering of exponent necessarily means an increase in the inequality of the events. Therefore, in that context, this current observation can provide an alternative to fitting the exponent values from a limited data set and just calculating their inequality could work in providing a potential hazard analysis.
In conclusion, we have studied earthqauke inequalities in models and real data for some earthqauke prone areas around the world and found consistent evidence that the earthqaukes become highly unequal prior to a large event. This can lead to potential earthqauke hazard analysis in future.
Acknowledgements
We thank B. K. Chakrabarti for fruitful discussions. The similations were performs in HPCC Chandrama at SRM University - AP.